### v1

# Faster Cascades via Speculative Decoding

Abstract
Cascades and speculative decoding are two common approaches to improving language models’ inference efficiency. Both approaches involve interleaving models of different sizes, but via fundamentally distinct mechanisms: cascades employ a deferral rule that invokes the larger model only for “hard” inputs, while speculative decoding uses speculative execution to primarily invoke the larger model in parallel verification mode. These mechanisms offer different benefits: empirically, cascades are often capable of yielding better quality than even the larger model, while theoretically, speculative decoding offers a guarantee of quality-neutrality. In this paper, we leverage the best of both these approaches by designing new speculative cascading techniques that implement their deferral rule through speculative execution. We characterize the optimal deferral rule for our speculative cascades, and employ a plug-in approximation to the optimal rule. Through experiments with T5 models on benchmark language tasks, we show that the proposed approach yields better cost-quality trade-offs than cascading and speculative decoding baselines.
1 Introduction
Large language models (LLMs) have demonstrated significant advances in quality on a range of natural language processing tasks [33, 34, 5, 2, 11, 48, 12, 41, 1, 43, 42], at the cost of a significant increase in inference latency. This has sparked a growing body of literature on reducing LMs’ inference costs without (overly) compromising on quality [15, 32, 36, 26, 7, 37, 40]. One such line of work involves constructing a family of models of various sizes (e.g., a small and large model), and suitably orchestrating amongst them to make a prediction. Two canonical instantiations of this strategy are model cascading [47, 28, 45, 23, 14, 8, 18, 13] and speculative decoding [39, 7, 26, 40, 49].
While similar in spirit, cascades and speculative decoding are fundamentally different in details. Cascades employ a deferral rule to identify “hard” inputs, and only invoke larger models on such inputs. For example, in a two-model cascade, one first invokes the smaller model, and uses its associated probability of the generated output to decide whether to defer to the larger model. By contrast, speculative decoding uses a small model to draft a block of tokens via standard auto-regressive decoding, which are then verified in parallel by a large model. One then accepts all drafted tokens until the first “implausible” one, which is rolled back based on the larger LM’s prediction.
Owing to their different mechnanisms, both methods have complementary strengths. Cascades seek to output distributions that have the best quality for a given cost budget, and are empirically observed to often yield better accuracies than even the individual models they are constructed with [22, 24]3). By contrast, speculative decoding is theoretically guaranteed to match the output distribution (or a close approximation thereof [44]), and are practically observed to provide impressive speed-ups [39, 7, 26, 40]. Given the complementary nature of these two approaches, a natural question that arises is whether we can leverage the best of both techniques.
In this paper, we do so by designing new techniques for two-model cascades that implement their deferral rule in a speculative manner: we have the smaller model generate drafts auto-regressively, and the larger model execute in parallel on the drafts to decide whether or not to defer on themWe show that this speculative cascading approach yields better cost-quality trade-offs than both standard cascades and speculative decoding. In detail, we make the following contributions:
  1. (i)
    We introduce a general recipe for speculative execution, where we seek to mimic a general target distribution that interleaves the drafter’s and verifier’s output distributions. Lossy speculative sampling [44] is a special case of this recipe for a particular target distribution (§4.1).
  2. (ii)
    We show how common cascading deferral rules, such as Chow’s rule [10] and confidence-difference thresholding [22], can be implemented speculatively by plugging in the corresponding target distribution into the proposed framework. We refer to these as speculative cascades4.2).
  3. (iii)
    We characterize the theoretically optimal deferral rule for a speculative cascade, and design a speculative cascading technique that implements a plug-in estimate to the optimal rule (Lem. 4).
  4. (iv)
    Through experiments on benchmark language tasks, we show that speculative cascades constructed from T5 models [35] of different sizes are able to provide better cost-quality trade-offs than their sequential cascade and speculative decoding counterparts (§6).
2 A Tale of Two Efficient LM Inference Strategies
Let $\mathscr{V}$ denote a finite vocabulary of tokens, with $\mathscr{V}^{*}$ the set of all sequences generated by this vocabulary. Let $\Delta_{\mathscr{V}}$ denote the set of all probability distributions over tokens in $\mathscr{V}$. Given an arbitrary length sequence $x=x_{1}x_{2}\ldots x_{L}\in\mathscr{V}^{*}$ and index $i\leq L$, denote by $x_{<i}=x_{1}x_{2}\ldots x_{i-1}$. A language model (LM) is a probability distribution over $\mathscr{V}^{*}$. Let $\mathbb{P}$ denote the ground-truth probability distribution over $\mathscr{V}^{*}$. This could be, for example, a distribution over prompt-response pairs that the LM may encounter during deployment, or a distribution of sequences used to pre-train the LM. We will measure the quality of an LM based on how closely it mimics $\mathbb{P}$.
Suppose we are provided two LMs $q$ and $p$, where $p$ is the larger model. Our goal is to design an inference strategy that selectively invokes $q$ and $p$ to trade-off between quality and latency (which may be approximated by the fraction of times that $p$ is invoked). We will denote by $q(x_{t}|x_{<t})$ the probability $q$ associates to token $x_{t}\in\mathscr{V}$ given prefix $x_{<t}\in\mathscr{V}^{t-1}$, and by $p(x_{t}|x_{<t})$ the same distribution from model $p$. Whenever it is clear from context, we will hide the conditioning on prefix $x_{<t}$, and use the shorthand $p_{t}(\cdot)$ for $p(\cdot|x_{<t})$ and $q_{t}(\cdot)$ for $q(\cdot|x_{<t})$.
Cascades.
Cascades are an effective strategy to trade-off cost and quality by having the smaller model $q$ handle the “easy” samples, and the larger model $p$ handle the “difficult” samples [18, 50]. A common approach to cascading is confidence-based thresholding or Chow’s rule [10, 22], where we first run $q$ on the given input, and defer to $p$ when $q$’s confidence for its generated response is sufficiently low. This strategy is typically implemented at the sequence-level, where for a given prefix $x_{<m}$ we invoke $q$ to generate a complete response $x_{m}\ldots x_{m+n}$. We evaluate the predicted probability from $q$ for this response, and check whether it falls below a threshold $\alpha\in[0,1]$:
$q(x_{m}\ldots x_{m+n}\,|\,x_{<m})<1-\alpha.$ (1)
If the above holds, we defer to $p$ to generate a new response; otherwise, we retain $q$’s response. One may then tune the threshold to achieve a desired cost-quality trade-off. The literature also offers variants of Chow’s rule that use a more nuanced aggregation of per-token uncertainties [18].
Speculative decoding.
Speculative decoding is an alternate inference strategy that applies token-level interleaving between $q$ and $p$, resulting in provably matching the larger model quality at a reduced inference cost [39, 26]. Given a prefix $x_{<t}$, we sample $\gamma$ draft tokens $x_{t},\ldots,x_{t+\gamma-1}$ auto-regressively from $q$, and run $p$ in parallel on the $\gamma$ prefixes $x_{<t},\ldots,x_{<t+\gamma-1}$, and verify if the generated tokens can be accepted. We then rollback to the first token $t+j^{*}$ that was rejected, replace $x_{t+j^{*}}$with a new token, and repeat the drafting-verification process with $x_{<t+j^{*}+1}$ as the new prefix.
During the verification stage, a draft token $x_{t+j}$ generated by $q$ is accepted with probability $\min\left(1,\frac{p_{t+j}(x_{t+j})}{q_{t+j}(x_{t+j})}\right)$ and rejected otherwise, where recall the shorthand $q_{t+j}(\cdot)=q(\cdot|x_{<t+j})$ and $p_{t+j}(\cdot)=p(\cdot|x_{<t+j})$. A rejected token is then replaced by a new token sampled from a modified distribution $\operatorname{norm}\left(\max\left\{0,\,p_{t+j}(\cdot)-q_{t+j}(\cdot)\right\}\right),$ where $\text{norm}(\cdot)$ denotes normalization to sum to 1. This sampling process is provably equivalent to sampling $\gamma$ tokens auto-regressively from $p$ for prefix $x_{<t}$ [26]. We summarize this speculative sampling procedure in Algorithm 1. Each invocation of this algorithm generates at most $\gamma+1$ next tokens for a given prefix $x_{<t}$. One may run this algorithm multiple times to generate a complete output sequence.
In practice, one may employ a lossy variant [44] of the above sampling that allows some deviation from verifier’s distribution $p$. In this case, a draft token $x_{t+j}$ is accepted with probability $\min\left(1,\frac{p_{t+j}(x_{t+j})}{(1-\alpha)\cdot q_{t+j}(x_{t+j})}\right)$, where $\alpha\in[0,1)$ is a strictness parameter, with higher values indicating greater deviation from $p$. A rejected token may then be replaced by a token sampled from the residual distribution $\operatorname{norm}\left(\max\left\{0,\,\frac{1}{\beta}\cdot p_{t+j}(\cdot)-q_{t+j}(\cdot)\right\}\right),$ where $\beta\geq 1-\alpha$ is a parameter that depends on $\alpha$, $q$ and $p$. A common heuristic is to simply set $\beta=1$ [52].
Inference strategy Deferral decision $\delta(q,p)$ Target distribution $\pi(x)$ Execution
SpecDecoding [26] - $p(x)$ Speculative
Lossy SpecDecoding [44] - $\max\{\min\{q(x),\frac{p(x)}{1-\alpha}\},\frac{p(x)}{\beta}\}$ Speculative
BiLD* [24] $\mathbf{1}\big{(}\,D(q,p)>\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
TokenCascade [Chow] [10] $\mathbf{1}\big{(}\max_{v}q(v)<1-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Sequential
Oracle [Diff] [22] $\mathbf{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Oracle
SpecCascade [Chow] $\mathbf{1}\big{(}\max_{v}q(v)<1-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [Diff] $\mathbf{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [OPT] $\mathbf{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\cdot D_{\textrm{TV}}(p,q)\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
Table 1: Target distributions associated with different inference algorithms, where $\alpha$ is a free parameter and $\beta\geq 1-\alpha$ is a parameter dependent on $\alpha$, $q$ and $p$. The last column indicates whether the execution is sequential (Algorithm 2), via an oracle (Algorithm 3), or speculative (Algorithm 5) [26]. The third row presents a variant of the BiLD algorithm of [24], where $D(q,p)$ is a measure of discrepancy between $q$ and $p$; the original algorithm differs in the use of a deterministic speculative decoding procedure with a dynamic draft window (see §5). See (6) for more details on $\delta$.
3 Cascades Meet Speculative Decoding
Both cascades and speculative decoding interleave models of different sizes to reduce inference cost, but fundamentally differ in the mechanisms they use. As a step towards comparing the strengths and weaknesses of these approaches, we first describe how one may design a token-level cascade.
3.1 Warm-up: Token-level cascades
It is straight-forward to extend the sequence-level Chow’s rule from §2 to form a token-level cascade between $q$ and $p$. For a prefix $x_{<t}$, we first compute the smaller model’s distribution $q(\cdot|x_{<t})$, and check whether $\max_{v\in\mathscr{V}}\,q(v|x_{<t})$ is below a pre-chosen threshold. if so, we evaluate $p(\cdot|x_{<t})$, and sample $x_{t}\sim p(\cdot|x_{<t})$; otherwise, we sample $x_{t}\sim q(\cdot|x_{<t})$.
More generally, we may design a token-level deferral rule ${r}:\mathscr{V}^{t-1}\rightarrow\{0,1\}$ that takes the prefix $x_{<t}$ as input and outputs a binary decision, with ${r}(x_{<t})=1$ indicating that we defer to $p$ (i.e., draw a sample from $p$ rather than $q$). For example, token-level Chow’s rule can be written as:
${r}_{\rm\tt Chow}(x_{<t})=1~{}\iff~{}\max_{v\in\mathscr{V}}\,q(v|x_{<t})<1-\alpha,$ (2)
where $\alpha$ is a cost parameter; the higher the value, the lower is the frequency of deferral to $p$. One may also use a confidence measure different from the maximum probability, such as the entropy of the small model’s probability distribution. We elaborate in §B that the choice of confidence measure would depend on the evaluation metric of interest; (2) is typically prescribed when the cascade’s quality is evaluated in terms of its accuracy against the ground-truth distribution on individual tokens, whereas entropy is prescribed when the metric of interest is the cross-entropy loss.
3.2 Optimal token-level cascade deferral
While Chow’s rule (2) is easy to implement, it can be sub-optimal if the smaller model’s maximum token probability is not reflective of which of the two models are better equipped to predict the next token for a given prefix [22]. Given this, it is natural to ask what the optimal deferral rule $r$ for a token-cascade looks like, and whether we can reasonably approximate this rule.
For this, we must first specify an objective to minimize at each step $t$. Following the prior cascade literature [22, 18], a reasonable objective to minimize is the expected loss from the deferral rule against the ground-truth distribution $\mathbb{P}$, with an added cost for deferring to the larger model. We state this below for a fixed prefix $x_{<t}$, using the short-hand $q_{t}(\cdot)$ for $q(\cdot|x_{<t})$ and $p_{t}(\cdot)$ for $p(\cdot|x_{<t})$:
$L_{\rm def}(r;x_{<t})$ $=\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\Big{[}\big{(}1-r(x_{<t})\big{)}\cdot\ell(x_{t},q_{t})+r(x_{<t})\cdot\big{(}\ell(x_{t},p_{t})+\alpha\big{)}\Big{]},$ (3)
for a cost penalty $\alpha\geq 0$ and loss function $\ell:\mathscr{V}\times\Delta_{\mathscr{V}}\rightarrow\mathbb{R}_{+}$. Common choices for $\ell$ include the 0-1 loss $\ell_{\text{0-1}}(x_{t},q_{t})=\mathbf{1}\left(x_{t}\neq\operatorname{argmax}_{v}q_{t}(v)\right)$ and the log loss $\ell_{\log}(x_{t},q_{t})=-\log\left(q_{t}(x_{t})\right).$
Lemma 1 (Optimal deferral for token-level cascades [22]).
The minimizer of (3) is of the form:
$r^{*}(x_{<t})=1~{}~{}\iff~{}~{}\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\,>\,\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]+\alpha.$ (4)
Intuitively, we compare the expected loss from $q$ with the expected cost of invoking $p$, and decide to defer when the latter is smaller. We note here that this optimization problem is set up for a fixed prefix $x_{<t}$. One may also consider the coupled optimization problem across all prefixes from 1 to $T$.
Plug-in estimator for (4). The optimal rule in (4) requires computing expectations over the ground-truth distribution $\mathbb{P}(\cdot|x_{>t})$, which is not available during inference time. A common approach in the cascades literature is to replace the expected losses with the models’ confidence estimates [22]. For example, when $\ell=\ell_{\text{0-1}}$, it may be reasonable to use $1-\max_{v}q_{t}(v)$ as an estimate of the expected 0-1 loss $\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell_{\text{0-1}}(x_{t},q_{t})\right]$ and $1-\max_{v}p_{t}(v)$ as an estimate of $\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell_{\text{0-1}}(x_{t},q_{t})\right]$. The extent to which these estimates are accurate depend on how well $q$ and $p$ are calibrated [17]. The resulting plug-in estimator for (4) is given by:
$\textstyle\hat{r}_{\rm\tt Diff}(x_{<t})=1~{}~{}\iff~{}~{}\max_{v}q_{t}(v)\,<\,\max_{v}p_{t}(v)-\alpha.$ (5)
Similarly, when $\ell=\ell_{\log}$, we may use the entropy $-\sum_{v}q_{t}(v)\cdot\log(q_{t}(v))$ from $q_{t}$ as an estimate of its expected log-loss, and similarly for $p_{t}$ (see §C).
Remark 1 (Oracle deferral rules).
$\hat{r}_{\rm\tt Diff}$ cannot be directly used in a token-level cascade, as it needs the larger model to be invoked at every position $t$. However, it serves as an oracle that allows to analyze the head-room available to improve upon Chow’s rule. See also Remark 2.
3.3 When do token-level cascades outperform speculative decoding?
Token-level cascades and speculative decoding differ in the distribution over tokens they seek to mimic. Speculative decoding seeks to mimic the larger model’s output distribution (or an approximation to it). As a result, the quality of the sampled output is limited by the accuracy of the larger model. On the other hand, token-level cascades seek to output distributions that closely approximate the ground-truth label distribution, and can potentially yield better quality than even the larger model.
Indeed, speculative decoding would be ideal in applications where the verification model $p$ is uniformly better than the draft model $q$ on all inputs. However, when there are inputs where the draft model fares better than the verifier, one may want to retain the drafter’s predictions even when it disagrees with the verifier. Similarly, in settings where both the drafter and verifier fare poorly on some inputs (e.g., due to label noise), one may again want to ignore the disagreement between the drafter and verifier, and avoid triggering unnecessary roll-backs.
As a concrete example, we consider token-level cascades of T5 models [35] of two different sizes finetuned on a WMT EN $\rightarrow$ DE translation [3] and an extreme summarization (XSum) task [31]. We construct these cascades using both (token-level) Chow’s rule in (2) and the oracle Diff rule in (5), and also apply speculative decoding with the smaller (larger) model as the drafter (verifier). In Figure 1, we plot quality as a function of fraction of samples deferred to the large model, as we vary the cost parameter $\alpha$. Note that with speculative decoding, each verification step is counted as a single call to the large model. While speculative decoding matches the quality of the large model (right-most point), the oracle rule yields significantly better quality on a wide range of operating points. Even Chow’s rule, which is sub-optimal for cascading [22], outperforms speculative decoding in a small region.
However, as also evident from the plots, token-level cascades require a significantly larger number of calls to the larger model to achieve the same quality. This is because token-level cascades are executed sequentially: whenever $q$ defers, we execute $p$ once to generate one next token for the prefix accumulated so far, and the control transfers back to $q$. In contrast, speculative decoding runs $p$ in scoring mode to verify multiple draft tokens from $q$ in parallel. Moreover, the stochastic verification algorithm in speculative decoding often results in fewer tokens from $q$ getting rejected compared to the deterministic deferral rules used in a cascade. These observations motivate a natural question: given their complementary strengths, how can we leverage the best of both these techniques?
Figure 1: Plot of quality as a function of deferral rate for cascades constructed from T5 models (under temperature sampling with $T=1$). The left-most point represents the small model and the right-most represents the large model. We compare token-level cascades constructed with Chow’s rule (Chow) and an oracle deferral rule (Diff), and speculative decoding with block size $\gamma=5$. We measure deferral rate in terms of the number of calls to the larger model; with a cascade, each call yields exactly one token, whereas with speculative decoding, a single call scores $\gamma$ draft tokens in parallel. While speculative decoding matches the quality of the large model (see dashed horizontal line), the oracle deferral rule yields significantly better quality on a range of deferral rates; this however comes at the cost of higher number of calls to the large model.
4 Speculative Cascades: Leveraging the Best of Both Worlds
In addressing the above question, we present our main contribution: a principled approach to combining the superior quality of cascades with the faster execution of speculative decoding approaches.
4.1 Speculative decoding with general target distributions
We begin by considering a generic version of speculative sampling that seeks to mimic a general target distribution derived from the drafter’s and verifier’s distributions. In the proposed sampling procedure outlined in Algorithm 4, we sample tokens auto-regressively as before from the drafter’s distribution. During the verification step, however, we do not compare the drafter’s token probabilities against the verifier’s distribution. Instead, we use a user-specified target distribution $\pi=\mathbb{T}(q,p)\in\Delta_{\mathscr{V}}$ derived from the drafter’s and verifier’s distributions at position $t$, for some function $\mathbb{T}(\cdot,\cdot)$ that is inexpensive to compute. We accept a draft token $x_{t}$ when $q(x_{t})\leq\pi(x_{t})$ and reject it otherwise with probability $1-\frac{\pi(x_{t})}{q(x_{t})}$. Upon rejection, we re-sample from the residual distribution $\operatorname{norm}\left(\max\{0,\pi(\cdot)-q(\cdot)\}\right)$.
This general procedure not only encompasses standard speculative decoding [26] for $\mathbb{T}(q,p)=p$, but also includes lossy speculative decoding [44] as a special case:
Lemma 2.
Algorithm 4 reduces to the lossy speculative sampling procedure in [44] with parameters $\alpha$ and $\beta$ when $\mathbb{T}(q,p)(v)=\max\{\min\{q(v),\frac{p(v)}{1-\alpha}\},\frac{p(v)}{\beta}\}$.
4.2 From sequential to speculative cascades
Equipped with Algorithm 4, we now propose new cascading techniques that implement their deferral rule in a speculative manner. Recall that a token-level cascade of two models $q$ and $p$ is defined by a deferral rule $r:\mathscr{V}^{t-1}\rightarrow\{0,1\}$. For a prefix $x_{<t}$, the next-token distribution at position $t$ modeled by this cascade can be written as:
$\pi(v)=(1-r(x_{<t}))\cdot q_{t}(v)+r(x_{<t})\cdot p_{t}(v).$
In fact, for all the deferral rules described in §2, the resulting distribution can be described by a target distribution function $\mathbb{T}_{\delta}$ of the form:
$\mathbb{T}_{\delta}(q,p)(v)=(1-\delta(q,p))\cdot q(v)+\delta(q,p)\cdot p(v),$ (6)
for some function $\delta:\Delta_{\mathscr{V}}\times\Delta_{\mathscr{V}}\rightarrow\{0,1\}$ that maps distributions $(q,p)$ to a binary decision. For example, for Chow, $\delta(q,p)=\mathbf{1}\big{(}\max_{v}q(v)<1-\alpha\big{)}$, and for Diff, $\delta(q,p)=\mathbf{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\big{)}.$ See Table 1 for a summary of target distributions for different deferral rules.
Our proposal is to then invoke the speculative sampling procedure in Algorithm 4 with $\mathbb{T}_{\delta}$ as the target distribution function. We outline this generic speculative cascading approach in Algorithm 5, and contrast it with the sequential execution of a deferral rule in Algorithm 2.
Remark 2 (Exact implementation of oracle deferral rule Diff).
Unlike a sequential cascade, where the larger model’s distribution $p$ is not available at the time the deferral decision is made (see Remark 1), with a speculative cascade, we can accommodate deferral rules like Diff that depend on both $q$ and $p$. This is because we run the larger model $p$ in parallel on drafts generated by the smaller model $q$, allowing us to compute both $p(\cdot)$ and $q(\cdot)$ on every prefix.
So far we have considered deferral rules designed for sequential cascades. In what follows, we derive the optimal deferral rule $r$ for a speculative cascade, where we sample speculatively from a target distribution $\pi=(1-r(x_{<t}))\cdot q_{t}+r(x_{<t})\cdot p_{t}$ using $q_{t}$ as the drafter.
Algorithm 1 SpecDecode
Models $q$, $p$, Prefix $x_{<t}$, Block size $\gamma$
$\mathbb{T}(q,p)\stackrel{{\scriptstyle.}}{{=}}p$
$\text{\tt{GenSpecSample}}(q,p,\mathbb{T},x_{<t},\gamma)$
Algorithm 2 TokenCascade
Models $q$, $p$, Deferral logic ${\delta}$, Prefix $x_{<t}$
$q_{t}(\cdot)\stackrel{{\scriptstyle.}}{{=}}q(\cdot|x_{<t})$
if ${\delta}(q_{t},\emptyset)=0$  then
    Sample $x_{t}\sim q_{t}(\cdot)$
else
    $p_{t}(\cdot)\stackrel{{\scriptstyle.}}{{=}}p(\cdot|x_{<t})$;   Sample $x_{t}\sim p_{t}(\cdot)$
end if
$x_{t}$
Algorithm 3 OracleCascade
Models $q$, $p$, Deferral logic ${\delta}$, Prefix $x_{<t}$
$q_{t}(\cdot)\stackrel{{\scriptstyle.}}{{=}}q(\cdot|x_{<t})$;    $p_{t}(\cdot)\stackrel{{\scriptstyle.}}{{=}}p(\cdot|x_{<t})$
if ${\delta}(q_{t},p_{t})=0$  then
    Sample $x_{t}\sim q_{t}(\cdot)$
else
    Sample $x_{t}\sim p_{t}(\cdot)$
end if
$x_{t}$
Algorithm 4 GenSpecSample
Models $q$, $p$, Target distr. $\mathbb{T}$, Prefix $x_{<t}$, Block size $\gamma$
$[\gamma]\equiv\{0,\ldots,\gamma\}$
Sample $\gamma$ tokens auto-regressively from $q$
for $j=0$ to $\gamma-1$ do
    $q_{t+j}(\cdot)\stackrel{{\scriptstyle.}}{{=}}q(\cdot|x_{<t+j});~{}~{}~{}~{}$ $x_{t+j}\sim q_{t+j}(\cdot)$
end for
Run $p$ in parallel to score $\gamma$ draft tokens
$p_{t+j}(\cdot)\stackrel{{\scriptstyle.}}{{=}}p(\cdot|x_{<t+j}),~{}\forall j\in[\gamma]$
$\pi_{t+j}=\mathbb{T}(q_{t+j},p_{t+j})$
Find the earliest draft token that gets rejected
$a_{j}\sim\text{Ber}\left(\min\left\{1,\frac{\pi_{t+j}({x_{t+j}})}{q_{t+j}({x_{t+j}})}\right\}\right),~{}\forall j\in[\gamma-1];$    $a_{\gamma}=0$
$j^{*}=\min\{j\in[\gamma]\,:\,a_{j}=0\}$
Sample a new token from residual distribution
$p_{\rm res}(\cdot)=\begin{cases}\operatorname{norm}\left(\max\left\{0,\,\pi_{t+j^{*}}(\cdot)-q_{t+j^{*}}(\cdot)\right\}\right)&\text{if}~{}j^{*}<\gamma\\ p_{t+\gamma}(\cdot)&\text{else}\end{cases}$
Sample $x_{t+j^{*}}\sim p_{\rm res}(\cdot)$
$x_{t},\ldots,x_{t+j^{*}}$
Algorithm 5 SpecCascade
Models $q$, $p$, Deferral logic $\delta$, Prefix $x_{<t}$, Block size $\gamma$
$\mathbb{T}_{\delta}(q,p)\stackrel{{\scriptstyle.}}{{=}}(1-\delta(q,p))\cdot q+\delta(q,p)\cdot p$
$\text{\tt{GenSpecSample}}(q,p,\mathbb{T}_{\delta},x_{<t},\gamma)$
4.3 Deferral risk for speculative cascades
As with sequential cascades (§2), we begin by defining an objective to minimize. We seek a deferral rule $r:\mathscr{V}^{t-1}\rightarrow\{0,1\}$ that minimizes a loss against the ground-truth distribution, while limiting the inference cost to be within a budget. (Per above, this deferral rule implicitly defines a target distribution $\pi$.) The inference cost crucially depends on how frequently a draft token is rejected in the verification phase, triggering a rollback. To this end, we derive the probability that a token sampled from $q$ is rejected during verification, for a target distribution resulting from a deferral rule $r$.
Lemma 3.
For a given prefix $x_{<t}$, and target distribution $\pi=(1-r(x_{<t}))\cdot q_{t}+r(x_{<t})\cdot p_{t}$, the probability of a token drawn from draft distribution $q_{t}$ being rejected is equal to:
$r(x_{<t})\cdot D_{\textup{{TV}}}(p_{t},q_{t}),$
where $D_{\textup{{TV}}}(p,q)=\sum_{v\in\mathscr{V}}\max\{0,p(v)-q(v)\}$ is the TV distance between $p$ and $q$.
Intuitively, whenever $r(x_{<t})=0$, $\pi(v)=q_{t}(v)$ and therefore there is no rejection or roll-back; when $r(x_{<t})=1$, the rejection rate equals $D_{\textup{{TV}}}(p_{t},q_{t})$.
For a fixed prefix $x_{<t}$, we formulate the goal of finding a solution to:
$\min_{r}$ $~{}~{}\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\Big{[}\big{(}1-r(x_{<t})\big{)}\cdot\ell(x_{t},q_{t})+r(x_{<t})\cdot\ell(x_{t},p_{t})\big{)}\Big{]}$
[…]
for some budget $B>0$. Equivalently, one may minimize an unconstrained objective similar to (3), for suitable cost parameter $\alpha>0$ (see §C.4):
$L_{\rm spec}(r;x_{<t})$ $=\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\Big{[}\big{(}1-r(x_{<t})\big{)}\cdot\ell(x_{t},q_{t})+r(x_{<t})\cdot\big{(}\ell(x_{t},p_{t})+\alpha\cdot D_{\textrm{TV}}(p_{t},q_{t})\big{)}\Big{]},$ (8)
Contrasting (8) with the deferral risk in (3) for a sequential cascade, a key difference is that the cost of deferring to the larger model is no longer a constant, but depends on the similarity between $q_{t}$ and $p_{t}$, as measured by the TV distance between them.
4.4 Optimal speculative deferral
We next derive the optimal deferral rule for (8), and construct a feasible estimator for it.
Lemma 4 (Optimal deferral for speculative cascades).
The minimizer of (8) is of the form:
$r^{*}(x_{<t})=1~{}~{}\iff~{}~{}\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\,>\,\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]+\alpha\cdot D_{\textup{{TV}}}(p_{t},q_{t}).$ (9)
When $p_{t}$ and $q_{t}$ are similar, the rejection rate for $q_{t}$ is low, and hence the deferral decision will depend largely on which of the two models yields a lower expected loss. When $p_{t}$ and $q_{t}$ are very different, the optimal decision is to defer to $p_{t}$ only when it yields a substantially lower loss than $q_{t}$.
Plug-in estimator for (9). The optimal rule requires estimating expectations with respect the ground-truth distribution $\mathbb{P}(\cdot|x_{<t}).$ We employ similar plug-in estimators as the ones used with sequential cascades (§3). When $\ell=\ell_{\text{0-1}}$, we replace the expected 0-1 loss with (one minus) the maximum probability from the model, giving us:
$\textstyle\hat{r}_{\rm\tt OPT}(x_{<t})=1~{}~{}\iff~{}~{}\max_{v}q_{t}(v)\,<\,\max_{v}p_{t}(v)-\alpha\cdot D_{\textup{{TV}}}(p_{t},q_{t}).$ (10)
The efficacy of the plug-in estimator depends on how closely the individual models approximate the ground-truth distribution $\mathbb{P}(\cdot|x_{<t})$; this is formalized by the following regret bound:
Lemma 5 (Regret bound for $\hat{r}_{\rm\tt OPT}$).
Suppose $\ell=\ell_{\emph{\text{0-1}}}$. Then for a fixed prefix $x_{<t}$:
$L_{\rm spec}(\hat{r}_{\rm\tt OPT};x_{<t})-\min_{r}\,L_{\rm spec}(r;x_{<t})~{}\leq~{}\max_{v\in\mathscr{V}}\big{|}\mathbb{P}(v|x_{<t})-q_{t}(v)\big{|}\,+\,\max_{v\in\mathscr{V}}\big{|}\mathbb{P}(v|x_{<t})-p_{t}(v)\big{|}.$
One can now run the speculative cascading procedure in Algorithm 5 using (10) as the deferral rule; the corresponding $\delta(\cdot)$ is listed in Table 1. See §C.2 for a similar derivation for $\ell=\ell_{\log}$.
Remark 3 (Optimal deferral with temperature sampling).
In practice it is common to apply a temperature scaling to both $q_{t}$ and $p_{t}$, and sample from $\tilde{q}_{t}(v)\propto q_{t}(v)^{1/T}$ and $\tilde{p}_{t}(v)\propto p_{t}(v)^{1/T}$ respectively, for a temperature parameter $T>0$. In this case, the constrained problem in (7) would use the TV distance between the temperature-scaled distributions $\tilde{q}_{t}$ and $\tilde{p}_{t}$ to measure the resulting rejection rate, and the optimal deferral rule in Lemma 4 would now use $D_{\textup{{TV}}}(\tilde{p}_{t},\tilde{q}_{t})$ instead of $D_{\textup{{TV}}}({p}_{t},{q}_{t})$. To construct a plug-in estimator to this optimal rule, we still prescribe using the unscaled probabilities $q_{t}$ and $p_{t}$ to estimate the expected loss, giving us, for $\ell=\ell_{\emph{\text{0-1}}}$:
$\textstyle\tilde{r}_{\rm\tt OPT}(x_{<t})=1~{}~{}\iff~{}~{}\max_{v}q_{t}(v)\,<\,\max_{v}p_{t}(v)-\alpha\cdot D_{\textup{{TV}}}(\tilde{p}_{t},\tilde{q}_{t}).$
One may run Algorithm 5 with $\tilde{r}_{\rm\tt OPT}$ as the deferral rule and the temperature scaled $\tilde{q}_{t}$ as the drafter.
Remark 4 (Special case of greedy decoding).
When temperature $T\rightarrow 0$, $D_{\textup{{TV}}}(\tilde{p}_{t},\tilde{q}_{t})=1$ whenever $\operatorname{argmax}_{v}p_{t}(v)\neq\operatorname{argmax}_{v}q_{t}(v)$, and is zero otherwise. In this case, running Algorithm 5 with $\tilde{r}_{\rm\tt OPT}$ as the deferral rule (and $\tilde{q}_{t}$ as the drafter) is equivalent to running it with $\hat{r}_{\rm\tt Diff}$ in (5) as the deferral rule. In other words, for greedy decoding, the optimal deferral rules for a speculative cascade coincides with that for a sequential cascade. We formalize this in Lemma 8 in §C.3.
5 Further related work
There has been a stream of work on improving the draft generation process in speculative decoding; these include having the drafter and verifier share the same backbone [39, 25, 6, 30, 21, 51, 16, 27], using multiple small draft models [9, 46], using tree-structured draft batches [38, 29], distilling the drafter with the verifier [52], and leveraging multiple sampled candidates from the drafter [40].
The work that is most closely related to our specific proposal is the Big Little Decoder (BiLD) [24], which can be seen as another lossy variant of speculative decoding [26, 44, 52]. BiLD has two phases: a fallback phase, during which the drafter $q$ is run auto-regressively until its maximum predicted probability is sufficiently low; and a rollback phase, during which the verifier $p$ is run in parallel on the prefixes generated by $q$ and rolls back to the point where $D(q,p)>\alpha$, for a metric $D$ that measures discrepancy and threshold $\alpha$. The fallback phase can be seen as implementing Chow’s deferral rule (2), and allows for the draft window size to vary dynamically based on an estimate of how likely the draft tokens will be accepted; the rollback phase can be seen as a deterministic variant of the rejection sampling algorithm of [26].
An advantage of BiLD over the rejection sampling algorithm in [26] is the use of Chow’s rule to vary the draft window size. However, the final target distribution it seeks to mimic, $\mathbb{T}_{\textup{BiLD}}(q,p)(v)=\mathbf{1}(D(q,p)\leq\alpha)\cdot q(v)+\mathbf{1}(D(q,p)>\alpha)\cdot p(v)$, is an approximation to $p$, in that the target distribution $\pi=\mathbb{T}_{\textup{BiLD}}(q,p)$ is chosen to satisfy $D(\pi,p)\leq\alpha$. Hence, in cases where $q$ deviates substantially from $p$, BiLD would choose $p$ as the target distribution, even $q$ offers better quality (as captured by a suitable loss function in §4.4) on a prefix. In contrast, our proposed approach in §4 uses speculative decoding to approximate target distributions that optimally cascade between $q$ and $p$. In our experiments, we compare the efficacy of using $\mathbb{T}_{\textup{BiLD}}$ as the target distribution with the target distributions we propose in this paper (see Table 1). propose in this paper (see Table 1).
6 Experimental results
We present experiments to demonstrate that the proposed speculative cascading techniques yield better cost-quality trade-offs compared to both sequential token-level cascades and standard speculative decoding. We construct speculative cascades with three different deferral rules: (i) Chow in (2), (ii) Diff in (5) and (iii) OPT in (10). Our experimental setup is based on [52]; see D.1 for details.
Tasks and models.
[…]
Baselines.
Since our primary goal is to demonstrate the benefits of combining both cascading and speculative decoding, we compare against representative methods from both paradigms:
  1. (i)
    Sequence-level cascade [22, 18] based on sequence-level Chow’s rule in (1) (SeqCascade [Chow])
  2. (ii)
    Token-level cascade outlined in Algorithm 2, with token-level Chow’s rule in (2) used for deferral [10, 19] (TokenCascade [Chow])
  3. (iii)
    Lossy speculative decoding described in §2, with both $\beta=1$ [26, 52] (SpecDecode [Lossy]) and with $\beta$ tuned using the procedure in [44] (SpecDecode [Lossy])
  4. (iv)
    A variant of the Big-Little Decoder approach [24], which applies Algorithm 4 to the target distribution $\mathbb{T}_{\textup{BiLD}}$ in §5 (BiLD).
We do not include the oracle approach in Algorithm 3 as it is impractical to evaluate its running time. We also note that, while one could potentially integrate other recent variants of the speculative decoding (e.g., those in §5 that warrant changes to the model architectures) into our speculative cascades framework, in the interest of a fair comparison, we use the original version by [26].
We evaluate all methods under both greedy decoding ($T=0$), and temperature sampling with temperature $T=1$. As noted in §4.3, with greedy decoding, the OPT deferral rule coincides with the Diff deferral rule. We set the block size $\gamma=5$ for the methods based on speculative execution.
Results.
In Figure 2, we present plots of quality versus latency for the different methods. In each case, we vary the lenience parameter $\alpha$, and plot either the BLEU or ROUGE-2 metric as a function of the relative latency to the larger model. For brevity, we include the three main baselines; in §D.3 we compare to SpecDecode [Lossy] [44].
Clearly, methods that use speculative execution are considerably faster than sequential token-level cascades, although sequential cascades do have a slight advantage in the low-latency regimes (unlike speculative approaches, which always call the large model after every $\gamma$ steps, a sequential cascade only invokes the large model when the small model defers). Under temperature sampling ($T=1$), the OPT speculative cascading strategy is seen to provide the best quality metrics for most latency values, with Diff coming in a close second. A similar trend is also seen with greedy decoding, where the optimal deferral strategy (SpecCascade [Diff]) often yields better quality trade-offs than the other methods, although the gap to speculative decoding is much smaller.
In Table 2, we report (i) the reduction in latency from different methods when matching the quality of the large model, and (ii) the best quality that each method can deliver without exceeding the latency of the large model (for temperature $T=1$). SpecCascade [OPT] is seen to yield the maximum speed-up and the best quality metrics. The cascading approaches are often seen to fare poorly on both quality and latency metrics, with the exception of WMT, where SeqCascade yields non-trivial speed-ups.
Table 2: Reduction in latency from different methods when matching the quality of the large model (cols 2–7), and the best quality metric when matching each method yields without exceeding the latency of the large model (cols 8–13). Quality is measured in terms of the BLEU score for WMT and ROUGE-2 for XSum and CNNDM. All results are under temperature sampling with $T=1$.
[…]
Small $\rightarrow$ Large Small $\rightarrow$ XL Small $\rightarrow$ Large Small $\rightarrow$ XL
Method WMT XSum CNNDM WMT XSum CNNDM WMT XSum CNNDM WMT XSum CNNDM
SeqCascade [Chow] 1.55$\times$ 0.84$\times$ 0.98$\times$ 2.46$\times$ 0.93$\times$ 0.94$\times$ 16.56 12.97 9.91 16.29 16.40 11.18
TokenCascade [Chow] 1.03$\times$ 0.93$\times$ 1.40$\times$ 1.46$\times$ 0.82$\times$ 1.51$\times$ 16.52 13.30 10.36 16.65 17.09 11.44
SpecDecode [Lossy] 1.61$\times$ 1.10$\times$ 1.57$\times$ 2.17$\times$ 1.28$\times$ 2.07$\times$ 17.26 13.90 10.43 16.94 17.36 11.53
BiLD 1.34$\times$ 1.04$\times$ 1.38$\times$ 1.85$\times$ 1.28$\times$ 1.84$\times$ 16.49 13.81 10.14 15.90 17.35 11.35
SpecCascade [Chow] 1.43$\times$ 1.04$\times$ 1.41$\times$ 2.01$\times$ 1.28$\times$ 1.97$\times$ 17.76 13.82 10.28 16.35 17.36 11.39
SpecCascade [Diff] 1.79$\times$ 1.17$\times$ 1.75$\times$ 2.44$\times$ 1.30$\times$ 2.15$\times$ 18.04 14.00 10.64 18.07 17.37 11.67
SpecCascade [OPT] 1.95$\times$ 1.17$\times$ 1.80$\times$ 2.61$\times$ 1.34$\times$ 2.21$\times$ 18.33 14.10 10.86 18.09 17.48 11.85
Figure 2: Plots of quality vs. latency. Each method interleaves T5-small with T5-large (or T5-XL). The $x$-axis tracks the latency relative to that of calling the large model on all inputs. The horizontal dotted line denotes the quality of the large model. See Appendix D for additional plots.
7 Conclusions
We have proposed new speculative cascading techniques that use a combination of auto-regressive drafting and parallel verification to implement their deferral rule, and shown that they yield better cost-quality trade-offs than standard cascades and speculative decoding. In the future, we wish to replace our plug-in estimators with a router model [18] trained explicitly on ground-truth samples to approximate the optimal rule. We also wish to improve the deferral objective we seek to optimize at each position $t$ (8), and replace it with a global (coupled) deferral objective that takes all prefixes from 1 to $T$ into account. Another useful direction to explore would be to extend our proposal to handle a general cascade with more than two models.
Appendix A Proofs
A.1 Proof of Lemma 1
Proof.
Expanding the loss in (3), we have:
$L_{\rm def}(r;x_{<t})$ $=\big{(}1-r(x_{<t})\big{)}\cdot\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]+r(x_{<t})\cdot\big{(}\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]+\alpha\big{)}$
$=r(x_{<t})\cdot\left(\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]+\alpha-\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\right)\,+\,\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]$
This objective is minimized by a deferral rule $r:\mathscr{V}^{t-1}\rightarrow\{0,1\}$ that minimizes, for each prefix $x_{<t}$, the term within the parenthesis. Therefore the minimizer $r^{*}(x_{<t})=1$ whenever the term within the parenthesis is negative:
$\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]+\alpha-\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]<0,$
and $r^{*}(x_{<t})=0$ otherwise. Re-arranging the terms completes the proof. ∎
A.2 Proof of Lemma 2
Proof.
The proof follows straight-forwardly from the results in [44]. Recall from §2 that the lossy speculative decoding procedure of [44] accepts a draft token $x$ with probability:
$\kappa(x)=\min\left\{1,\frac{p(x)}{(1-\alpha)\cdot q(x)}\right\},$ (11)
and replaces a rejected draft token with a token sampled from the residual distribution:
$p_{\rm res}(x)=\operatorname{norm}\left(\max\left\{0,\,\frac{1}{\beta}\cdot p(x)-q(x)\right\}\right),$ (12)
for parameters $\alpha\in[0,1)$ and $\beta\geq 1-\alpha$.
We need to show that running Algorithm 4 with the target distribution:
$\pi(x)=\max\left\{\min\left\{q(x),\frac{p(x)}{1-\alpha}\right\},\frac{p(x)}{\beta}\right\}$
results in the same acceptance probability (11) and residual distribution (12).
The acceptance probability for a draft token $x$ when running Algorithm 4 on $\pi$ is given by:
$\kappa^{\pi}(x)=\min\left\{1,\frac{\pi(x)}{q(x)}\right\}.$
The corresponding residual distribution is given by:
$p^{\pi}_{\rm res}(x)=\operatorname{norm}\left(\max\left\{0,\,\pi(x)-q(x)\right\}\right).$
We consider three possible cases:
Case (i): $q(x)>\frac{1}{1-\alpha}\cdot p(x)\geq\frac{1}{\beta}\cdot p(x)$. In this case, $\pi(x)=\frac{1}{1-\alpha}\cdot p(x)$. As a result:
$\kappa^{\pi}(x)=\min\left\{1,\frac{p(x)}{(1-\alpha)\cdot q(x)}\right\}=\kappa(x);$
$p^{\pi}_{\rm res}(x)$ $=\operatorname{norm}\left(\max\left\{0,\,\frac{1}{1-\alpha}\cdot p(x)-q(x)\right\}\right)$
$=0=\operatorname{norm}\left(\max\left\{0,\,\frac{1}{\beta}\cdot p(x)-q(x)\right\}\right)=p_{\rm res}(x).$
Case (ii): $\frac{1}{1-\alpha}\cdot p(x)\geq\frac{1}{\beta}\cdot p(x)>q(x)$. In this case, $\pi(x)=\frac{1}{\beta}\cdot p(x)$. As a result:
$\kappa^{\pi}(x)=\min\left\{1,\frac{p(x)}{\beta\cdot q(x)}\right\}=1=\min\left\{1,\frac{p(x)}{(1-\alpha)\cdot q(x)}\right\}=\kappa(x);$
$p^{\pi}_{\rm res}(x)$ $=\operatorname{norm}\left(\max\left\{0,\,\frac{1}{\beta}\cdot p(x)-q(x)\right\}\right)=p_{\rm res}(x).$
Case (iii): $\frac{1}{1-\alpha}\cdot p(x)\geq q(x)\geq\frac{1}{\beta}\cdot p(x)$. In this case, $\pi(x)=q(x)$. As a result:
$\kappa^{\pi}(x)=1=\min\left\{1,\frac{p(x)}{(1-\alpha)\cdot q(x)}\right\}=\kappa(x);$
$p^{\pi}_{\rm res}(x)=0=\operatorname{norm}\left(\max\left\{0,\,\frac{1}{\beta}\cdot p(x)-q(x)\right\}\right)=p_{\rm res}(x).$
In all three cases, the acceptance probabilities and residual distributions are identical. ∎
A.3 Proof of Lemma 3
Proof.
Under a target distribution $\pi_{t}$, the probability of a draft token drawn from $q_{t}$ being is rejected is given by [26]:
rejection probability $=\sum_{v\in\mathscr{V}}q_{t}(v)\cdot\left(1-\min\left\{1,\frac{\pi_{t}(v)}{q_{t}(v)}\right\}\right)$
$=1-\sum_{v\in\mathscr{V}}\min\left\{q_{t}(v),\pi_{t}(v)\right\}$
$=\sum_{v\in\mathscr{V}}\pi_{t}(v)-\sum_{v\in\mathscr{V}}\min\left\{q_{t}(v),\pi_{t}(v)\right\}$
$=\sum_{v\in\mathscr{V}}\max\left\{0,\pi_{t}(v)-q_{t}(v)\right\}.$
Expanding $\pi$, the rejection probability becomes:
rejection $\text{rejection probability}=\sum_{v\in\mathscr{V}}\max\left\{0,\,(1-r(x_{<t}))\cdot q_{t}(v)+r(x_{<t})\cdot p_{t}(v)-q_{t}(v)\right\}$
When $r(x_{<t})=1$, we have:
rejection probability $=\sum_{v\in\mathscr{V}}\min\left\{0,\,p_{t}(v)-q_{t}(v)\right\}=D_{\textrm{TV}}(p_{t},q_{t})=r(x_{<t})\cdot D_{\textrm{TV}}(p_{t},q_{t}).$
When $r(x_{<t})=0$, we have:
rejection probability $=0=r(x_{<t})\cdot D_{\textrm{TV}}(p_{t},q_{t}),$
as desired. ∎
A.4 Proof of Lemma 4
Proof.
Expanding the deferral risk in (8), we have:
$L_{\rm spec}(r;x_{<t})$ $=r(x_{<t})\cdot\left(\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]+\alpha\cdot D_{\textrm{TV}}(p_{t},q_{t})-\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\right)$
$\hskip 199.16928pt+\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right].$
This objective is minimized by a deferral rule $r:\mathscr{V}^{t-1}\rightarrow\{0,1\}$ that minimizes, for each prefix $x_{<t}$, the term within the parenthesis. Therefore the minimizer $r^{*}(x_{<t})=1$ whenever the term within the parenthesis is negative:
$\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]+\alpha\cdot D_{\textrm{TV}}(p_{t},q_{t})-\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]<0,$
and $r^{*}(x_{<t})=0$ otherwise. Re-arranging the terms completes the proof. ∎
A.5 Proof of Lemma 5
For a fixed prefix $x_{<t}$, we can write the deferral risk in (8) as:
$L_{\rm spec}(r;x_{<t})$ $=r(x_{<t})\cdot\left(\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]+\alpha\cdot D_{\textrm{TV}}(p_{t},q_{t})-\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\right)+C,$
where $C$ is a term independent of the deferral rule $r$. Let $r^{*}:\mathscr{V}^{t-1}\rightarrow\{0,1\}$ denote the optimal deferral rule that minimizes $L_{\rm spec}$ for any prefix $x_{<t}$. We then have:
$L_{\rm spec}\left(\hat{r}_{\rm\tt OPT};x_{<t}\right)-L_{\rm spec}\left(r^{*};x_{<t}\right)$
$=\left(\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right)\cdot\left(\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]+\alpha\cdot D_{\textrm{TV}}(p_{t},q_{t})-\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\right).$
Adding and subtracting $\max_{v}q_{t}(v)-\max_{v}p_{t}(v)$ to the term within the second parenthesis, we get:
$L_{\rm spec}\left(\hat{r}_{\rm\tt OPT};x_{<t}\right)-L_{\rm spec}\left(r^{*};x_{<t}\right)$
$=\left(\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right)\cdot\left(\max_{v}q_{t}(v)+\alpha\cdot D_{\textrm{TV}}(p_{t},q_{t})-\max_{v}p_{t}(v)\right)$
$~{}~{}~{}~{}~{}+\left(\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right)\cdot\left(\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]-\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]-\max_{v}q_{t}(v)+\max_{v}p_{t}(v)\right)$
$=\left(\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right)\cdot\left(\max_{v}p_{t}(v)+\alpha\cdot D_{\textrm{TV}}(p_{t},q_{t})-\max_{v}q_{t}(v)\right)$
$~{}~{}~{}~{}~{}+\left(\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right)\cdot\left(\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]-1+\max_{v}p_{t}(v)\right)$
$~{}~{}~{}~{}~{}+\left(\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right)\cdot\left(1-\max_{v}q_{t}(v)-\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\right)$
$=\left(\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right)\cdot\left(\max_{v}p_{t}(v)+\alpha\cdot D_{\textrm{TV}}(p_{t},q_{t})-\max_{v}q_{t}(v)\right)$
$~{}~{}~{}~{}~{}+\left|\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right|\cdot\left|\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]-1+\max_{v}p_{t}(v)\right|$
$~{}~{}~{}~{}~{}+\left|\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right|\cdot\left|1-\max_{v}q_{t}(v)-\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\right|$
$=\underbrace{\left(\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right)\cdot\left(\max_{v}p_{t}(v)+\alpha\cdot D_{\textrm{TV}}(p_{t},q_{t})-\max_{v}q_{t}(v)\right)}_{\text{term}_{1}}$
$~{}~{}~{}~{}~{}+\underbrace{\left|\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]-1+\max_{v}p_{t}(v)\right|}_{\text{term}_{2}}+\underbrace{\left|1-\max_{v}q_{t}(v)-\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\right|}_{\text{term}_{3}}$ (13)
where we have used the fact that $\left|\hat{r}_{\rm\tt OPT}(x_{<t})-r^{*}(x_{<t})\right|\leq 1.$
We bound each term separately. For the first term, consider two cases: (i) $\max_{v}p_{t}(v)+\alpha\cdot D_{\textrm{TV}}(p_{t},q_{t})-\max_{v}q_{t}(v)\leq 0$ and (ii) $\max_{v}p_{t}(v)+\alpha\cdot D_{\textrm{TV}}(p_{t},q_{t})-\max_{v}q_{t}(v)>0$. When (i) holds, $\hat{r}_{\rm\tt OPT}(x_{<t})=1$; so irrespective of whether $r^{*}(x_{<t})$ is 0 or 1,
$\text{term}_{1}$ $\leq\max_{v}p_{t}(v)+\alpha\cdot D_{\textrm{TV}}(p_{t},q_{t})-\max_{v}q_{t}(v)\leq 0$
When (ii) holds, $\hat{r}_{\rm\tt OPT}(x_{<t})=0$; so irrespective of whether $r^{*}(x_{<t})$ is 0 or 1,
$\text{term}_{1}$ $\leq-\left(\max_{v}p_{t}(v)+\alpha\cdot D_{\textrm{TV}}(p_{t},q_{t})-\max_{v}q_{t}(v)\right)<0.$
Thus we have:
$\text{term}_{1}\leq 0.$ (14)
We next move to the second term. Since $\ell=\ell_{\text{0-1}}$, we have:
$\text{term}_{2}$ $=\left|\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]-1+\max_{v}p_{t}(v)\right|$
$=\left|\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[1\left(x_{t}\neq\operatorname{argmax}_{v}p_{t}(v)\right)\right]-1+\max_{v}p_{t}(v)\right|$
$=\left|\max_{v}p_{t}(v)-\sum_{x_{t}}\mathbb{P}(x_{t}|x_{<t})\cdot 1\left(x_{t}=\operatorname{argmax}_{v}p_{t}(v)\right)\right|$
Suppose $v^{*}\in\operatorname{argmax}_{v}p_{t}(v)$, then:
$\text{term}_{2}$ $=\left|p_{t}(v^{*})-\mathbb{P}(v^{*}|x_{<t})\right|\leq\max_{v}\left|p_{t}(v)-\mathbb{P}(v|x_{<t})\right|.$ (15)
Similarly, we can show that:
$\text{term}_{3}$ $\leq\max_{v}\left|q_{t}(v)-\mathbb{P}(v|x_{<t})\right|.$ (16)
Substituting (14)–(16) in (13) completes the proof.
Inference strategy Deferral decision $\delta(q,p)$ Target distribution $\pi(x)$ Execution
SpecDecoding [26] - $p(x)$ Speculative
Lossy SpecDecoding [44] - $\max\{\min\{p(x),\frac{q(x)}{1-\alpha}\},\frac{q(x)}{\beta}\}$ Speculative
BiLD* [24] $\mathbf{1}\big{(}\,D(q,p)>\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
Cascade [Chow] [10] $\mathbf{1}\big{(}\max_{v}q(v)<1-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Sequential
Cascade [ChowLog] $\mathbf{1}\big{(}\,\textrm{entropy}(q)>\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Sequential
Oracle [Diff] [22] $\mathbf{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Oracle
Oracle [DiffLog] $\mathbf{1}\big{(}\,\textrm{entropy}(p)<\textrm{entropy}(q)-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Oracle
SpecCascade [Chow] $\mathbf{1}\big{(}\max_{v}q(v)<1-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [ChowLog] $\mathbf{1}\big{(}\,\textrm{entropy}(q)>\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [Diff01] $\mathbf{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [DiffLog] $\mathbf{1}\big{(}\,\textrm{entropy}(p)<\textrm{entropy}(q)-\alpha\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [OPT01] $\mathbf{1}\big{(}\max_{v}q(v)<\max_{v}p(v)-\alpha\cdot D_{\textrm{TV}}(p,q)\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
SpecCascade [OPTLog] $\mathbf{1}\big{(}\,\textrm{entropy}(p)<\textrm{entropy}(q)-\alpha\cdot D_{\textrm{TV}}(p,q)\big{)}$ $(1-\delta)\cdot q(x)+\delta\cdot p(x)$ Speculative
Table 3: Target distributions associated with different inference algorithms, where $\alpha$ is a free parameter and $\beta\geq 1-\alpha$ is a parameter dependent on $q,p$ and $\alpha$. The last column indicates whether the execution is sequential (Algorithm 2), via an oracle (Algorithm 3), or speculative (Algorithm 5) [26]. The third row presents a variant of the BiLD algorithm of [24], where $D(q,p)$ is a measure of discrepancy between $q$ and $p$; the original algorithm differs from [26] in the use of a deterministic speculative decoding procedure with a dynamic draft window (see §5).
Appendix B Derivation of Chow’s rule
We show below that Chow’s rule is a plug-in estimator to the optimal solution to the following objective
$L_{\rm rej}(r;x_{<t})$ $=\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\Big{[}\big{(}1-r(x_{<t})\big{)}\cdot\ell(x_{t},q_{t})+r(x_{<t})\cdot\alpha\Big{]},$ (17)
where the deferral rule is penalized with a constant penalty $\alpha\in[0,1]$ for choosing to defer to the large model.
Following the same steps as Lemma 1, it is easy to show:
Lemma 6.
The minimizer of (17) is of the form:
$r^{*}(x_{<t})=1~{}~{}\iff~{}~{}\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\,>\,\alpha.$ (18)
If $\ell=\ell_{\text{0-1}}$, one may employ a plug-in estimator to (18) by replacing the expected 0-1 loss over $q_{t}$ with $1-\max_{v}q_{t}(v)$, giving us $\hat{r}_{\rm\tt Chow}(x_{<t})$ in (2). If $\ell=\ell_{\log}$, one may replace the expected log loss over $q_{t}$ with the entropy of $q_{t}$, giving us:
$\hat{r}_{\rm\tt ChowLog}(x_{<t})=1~{}\iff~{}\textrm{entropy}\big{(}q(\cdot|x_{<t})\big{)}>\alpha,$ (19)
where $\textrm{entropy}(q)=-\sum_{v\in\mathscr{V}}q(v)\cdot\log(q(v)).$
Appendix C Optimal Deferral: Additional Discussion
We provide additional discussion for the optimal deferral rules derived in §3 and §4.3.
C.1 Optimal sequential deferral when $\ell=\ell_{\log}$
Recall that the optimal deferral rule for a sequential cascade in Lemma 1 takes the form:
$r^{*}(x_{<t})=1~{}~{}\iff~{}~{}\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\,>\,\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]+\alpha\cdot D_{\textup{{TV}}}(p_{t},q_{t}).$
When $\ell=\ell_{\log}$, we may use the entropy $-\sum_{v}q_{t}(v)\cdot\log(q_{t}(v))$ from $q_{t}$ as an estimate of its expected log-loss, and similarly for $p_{t}$, giving us the plug-in estimator:
$\textstyle\hat{r}_{\rm\tt DiffLog}(x_{<t})=1~{}~{}\iff~{}~{}\sum_{v}q_{t}(v)\cdot\log(q_{t}(v))\,<\,\sum_{v}p_{t}(v)\cdot\log(p_{t}(v))-\alpha.$ (20)
C.2 Optimal speculative deferral when $\ell=\ell_{\log}$
Recall that the optimal deferral rule for a speculative cascade in Lemma 4 takes the form:
$r^{*}(x_{<t})=1~{}~{}\iff~{}~{}\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},q_{t})\right]\,>\,\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\ell(x_{t},p_{t})\right]+\alpha\cdot D_{\textup{{TV}}}(p_{t},q_{t}).$
When $\ell=\ell_{\log}$, one may construct a plug-in estimator for the above rule by replacing the expected log loss with the entropy from the distribution:
$\textstyle\hat{r}_{\rm\tt OPTLog}(x_{<t})=1~{}\iff~{}\sum_{v}q_{t}(v)\cdot\log(q_{t}(v))\,<\,\sum_{v}p_{t}(v)\cdot\log(p_{t}(v))-\alpha\cdot D_{\textup{{TV}}}(p_{t},q_{t}).$ (21)
Lemma 7 (Regret bound for $\hat{r}_{\rm\tt OPTLog}$).
Suppose $\ell=\ell_{\log}$. Suppose for a fixed $x_{<t}$, $|\log(q(v))|\leq B_{q}$ and $|\log(p(v))|\leq B_{p},\,\forall v\in\mathscr{V}$, for some $B_{q},B_{p}>0$. Then:
$L_{\rm spec}(r_{\rm\tt OPT};x_{<t})-\min_{r}L_{\rm spec}(r;x_{<t})~{}\leq~{}\textstyle B_{q}\cdot\sum_{v\in\mathscr{V}}\big{|}\mathbb{P}(v|x_{<t})-q_{t}(v)\big{|}\,+\,B_{p}\cdot\sum_{v\in\mathscr{V}}\big{|}\mathbb{P}(v|x_{<t})-p_{t}(v)\big{|}.$
Proof.
The proof follows similar steps to that for Lemma 5, except in bounding the resulting $\text{term}_{2}$ and $\text{term}_{3}$ for the log loss. In this case,
$\text{term}_{2}$ $=\left|\mathbb{E}_{x_{t}\sim\mathbb{P}(\cdot|x_{<t})}\left[\log(p_{t}(x_{t}))\right]-\sum_{v}p_{t}(v)\cdot\log(p_{t}(v))\right|$
$=\left|\sum_{v}\mathbb{P}(v|x_{<t})\cdot\log(p_{t}(v))-\sum_{v}p_{t}(v)\cdot\log(p_{t}(v))\right|$
$\leq\sum_{v}\left|\mathbb{P}(v|x_{<t})-\sum_{v}p_{t}(v)\right|\cdot\log(p_{t}(v))$
$\leq B_{p}\cdot\sum_{v}\left|\mathbb{P}(v|x_{<t})-\sum_{v}p_{t}(v)\right|.$
Similarly,
$\text{term}_{3}$ $\leq\sum_{v}\left|\mathbb{P}(v|x_{<t})-\sum_{v}p_{t}(v)\right|\cdot\log(p_{t}(v))$
$\leq B_{q}\cdot\sum_{v}\left|\mathbb{P}(v|x_{<t})-\sum_{v}q_{t}(v)\right|.$
Plugging these bounds into the equivalent of (13) in Lemma 5 for the log-loss completes the proof. ∎
C.3 Optimal speculative deferral for greedy decoding
Lemma 8.
When $T\rightarrow 0$, running Algorithm 5 with $\tilde{r}_{\rm\tt OPT}$ as the deferral rule and $\tilde{q}_{t}$ as the drafter is equivalent to running it with $\hat{r}_{\rm\tt Diff}$ in (5) as the deferral rule and $\tilde{q}_{t}$ as the drafter.
Proof.
Note that under greedy inference, $\tilde{q}_{t}$ $\tilde{p}_{t}$ are one-hot encodings of $\operatorname{argmax}_{v}q_{t}(v)$ and $\operatorname{argmax}_{v}p_{t}(v)$ respectively. As a result,
$D_{\textrm{TV}}(\tilde{q}_{t},\tilde{p}_{t})=\mathbf{1}\left(\operatorname{argmax}_{v}q_{t}(v)\neq\operatorname{argmax}_{v}p_{t}(v)\right).$
When running Algorithm 5 with $\tilde{r}_{\rm\tt OPT}$ as the deferral rule, we will accept a draft token $v$ with probability:
$\kappa(v)=\min\left\{1,\frac{(1-\delta_{\rm OPT}(q,p))\cdot\tilde{q}(v)+\delta_{\rm OPT}(q,p)\cdot\tilde{p}(v)}{\tilde{q}(v)}\right\}$
where $\delta_{\rm OPT}(q,p)~{}=~{}\mathbf{1}\left(\max_{v}q(v)\,<\,\max_{v}p(v)-\alpha\cdot\mathbf{1}\left(\operatorname{argmax}_{v}q(v)\neq\operatorname{argmax}_{v}p(v)\right)\right)$. When $\operatorname{argmax}_{v}q(v)=\operatorname{argmax}_{v}p(v)$, then $\tilde{q}=\tilde{p}$, and irrespective of the outcome of $\delta(q,p),$ we have that $\pi(v)=1$. When $\operatorname{argmax}_{v}q(v)\neq\operatorname{argmax}_{v}p(v)$, then
$\pi(v)=1-\delta_{\rm OPT}(q,p)=\mathbf{1}\left(\max_{v}q(v)\,\geq\,\max_{v}p(v)-\alpha\right)=1-\delta_{\rm\tt Diff}(q,p).$
When a token gets rejected, we sample a new token from the residual distribution:
$p_{\rm res}(v)\propto\min\{0,(1-\delta_{\rm OPT}(q,p))\cdot\tilde{q}(v)+\delta_{\rm OPT}(q,p)\cdot\tilde{p}(v)-\tilde{q}(v)\}=\delta_{\rm OPT}(q,p)\cdot\min\{0,\tilde{p}(v)-\tilde{q}(v)\}$
When $\operatorname{argmax}_{v}q(v)=\operatorname{argmax}_{v}p(v)$, $p_{\rm res}(v)=0$. When $\operatorname{argmax}_{v}q(v)\neq\operatorname{argmax}_{v}p(v)$,
$p_{\rm res}(v)\propto\delta_{\rm OPT}(q,p)\cdot\min\{0,\tilde{p}(v)-\tilde{q}(v)\}=\delta_{\rm Diff}(q,p)\cdot\min\{0,\tilde{p}(v)-\tilde{q}(v)\}.$
Thus both the acceptance probability and the residual distribution are the same as the one we would have used had we run Algorithm 5 with $\hat{r}_{\rm\tt Diff}$ as the deferral rule. ∎
C.4 Equivalence between (7) and (8)
Since the prefix $x_{<t}$ is fixed in (7), the constrained optimization we seek to solve is of essentially of the following form:
$\min_{r\in\{0,1\}}(1-r)\cdot c_{0}+r\cdot c_{1}~{}~{}~{}\text{s.t.}~{}~{}~{}r\cdot c_{2}\leq B,$
for some coefficients $c_{0},c_{1},c_{2}>0$. Since $r$ is a binary variable, we may formulate an equivalent unconstrained problem with the same minimizer:
$\min_{r\in\{0,1\}}(1-r)\cdot c_{0}+r\cdot c_{1}+\alpha\cdot r\cdot c_{2},$
where we choose $\alpha=0$ when $c_{2}\leq B$ and choose an $\alpha>\frac{1}{c_{2}}\cdot(c_{0}-c_{1})$. This unconstrained optimization problem is of the form in (8).
Appendix D Additional Experimental Details
We provide additional details about our experimental setup and additional experimental results.
Figure 3: Plots of quality vs. latency. Each method interleaves T5-small with T-large (or T5-XL). The $x$-axis tracks the latency relative to that of calling the large model on all inputs. The horizontal dotted line denotes the quality of the large model. All results use temperature sampling with $T=1$.
Figure 4: Plots of quality vs. latency. Each method interleaves T5-small with T-large (or T5-XL). The $x$-axis tracks the latency relative to that of calling the large model on all inputs. The horizontal dotted line denotes the quality of the large model. All results use greedy decoding ($T=0$)
Figure 5: Plots of quality vs. latency comparing lossy speculative decoding with $\beta=1$ [Lossy] against $\beta$ tuned using the procedure in [44] [Lossy].
D.1 Experimental setup and hyper-parameters
We elaborate on our experimental setup and the hyper-parameters used.
Datasets. For the WMT English to German translation task [4], we use a validation sample of size 3000 provided with the dataset. We set the maximum input length to 80 and the maximum output length to 80. For the Extreme Summarization (XSum) task [31], we use a validation sample of size 11305, and set the maximum input length to 1024 and the maximum output length to 64. For the CNN/Daily Mail summarization task [20], we use a validation sample of size 13368, and set the maximum input length to 2048 and the maximum output length to 128. Following [52], we use ROUGE-2 as the evaluation metric for the summarization tasks.
Models. We construct cascades from T5 v1.1 family of encoder-decoder models [34], of different sizes T5-small (77M), T5-base (250M), T5-large (800M) and T5-XL (3B). [^1] The pre-trained checkpoints we use are available at: https://console.cloud.google.com/storage/browser/t5-data/pretrained_models. We follow the protocol in [52]: we initialize with the public checkpoints, pre-train them further for 100K steps, and supervise finetune pre-trained models on the three respective tasks. We finetune them for a maximum of 250K steps on WMT, a maximum of 100K steps on XSum and a maximum of 200K steps on CNNDM.
Evaluation. For each dataset, we evaluate the quality metrics on the entire validation set. For the run-time analysis, we adopt the protocol followed in [26, 52]. We randomly sample 500 examples from the validation set, and calculate the wall-clock time taken for decoding with a batch size of 1. We repeat this for three trials and report the average running time. All methods are run on the same TPUv4 device. The drafter and verifier models are run without model parallelism.
Hyper-parameters. We set the block-size $\gamma$ to 5 for all methods that use speculative execution. For the token-level cascades, we allow the small model to predict for a maximum of 10 tokens (similar to [24]), before invoking the large model. This was needed, as otherwise, the small model would predict a long sequence, and when it eventually defers to the large model, the large model is bottle-necked by the pre-filling of the long prefix accumulated by the small model. We vary the lenience parameter $\alpha$ to vary the latency and plot quality as a function of latency. We vary this parameter in the range 0 to 1 for all methods where the thresholding is on a probability metric; the exceptions to this are the BiLD variants, for which, we use a longer range, as detailed below.
BiLD baseline. For the BiLD method, we adopt the same discrepancy metric $D$ as [24] for greedy decoding:
$D(q,p)=\log\left(p\left(\operatorname{argmax}_{v}q(v)\right)\right),$
and pick the value of the threshold $\alpha$ on this metric from the range $[0,10]$. For temperature sampling with a non-zero temperature, we use the following natural analogue to the above $D$:
$D(q,p)=\mathbb{E}_{v\sim q}\left[\log(p(v))\right]=\sum_{v\in\mathscr{V}}q(v)\cdot\log(p(v)).$
Lossy speculative decoding. See §D.3 for details.
D.2 Additional experimental plots
In Figures 3 and 4, we provide additional plots of quality vs. latency for different inference strategies under temperature sampling ($T=1$) and greedy decoding respectively. Greedy decoding is seen to perform better compared to temperature sampling on the particular quality metrics we care about.
D.3 Lossy speculative decoding variants
In §6, we compared against the lossy speculative decoding [44, 52] described in §2, with the parameter $\beta$ set to 1. We now present results for this method with $\beta$ tuned according to the procedure in [44], and show that choosing $\beta=1$ fares at least as well as tuning $\beta$.
The goal in [44] is to choose $\alpha$ and $\beta$ so as to maximize the acceptance rate for the draft token, while ensuring that the KL divergence between the resulting target distribution and $p$ is within an allowable limit $R$. The authors prescribe specifying $R$, and for each prefix, tuning $\alpha$ and $\beta$ to solve the resulting constrained optimization problem. To be consistent with the rest of our experimental setup, we vary $\alpha$ to vary the draft acceptance rate (note that each choice of $\alpha$ corresponds to a particular KL divergence to $p$), and tune $\beta\geq 1-\alpha$ to satisfy the following condition outlined in [44]:
$\sum_{v}\max\left\{0,q(v)-\frac{p(v)}{1-\alpha}\right\}=\sum_{v}\max\left\{0,\frac{p(v)}{\beta}-q(v)\right\}$
We pick $\beta$ using a grid-search over 1000 values between $\alpha$ and 10. Since this tuning procedure, in turn, can add to the method’s latency, for a fair comparison, we analyze quality as a function of the fraction of calls to the large model. In Figure 5 (bottom), we plot these trade-off curves for loss speculative decoding with $\beta=1$ (Lossy) and for speculative decoding with $\beta$ tuned using the above procedure (Lossy). We compare performances on WMT and XSum, and in each case, interleave a T5-small model with a T5-large model.
In both cases, setting $\beta=1$ provides trade-offs comparable to or better than using a tuned value of $\beta$. The reason using a tuned value of $\beta$ fares worse than setting $\beta=1$ might be because we are measuring quality in terms of BLEU or ROUGE-2, which is different from the KL divergence to $p$ objective that the tuning procedure in [44] seeks to optimize.